OpenStudy (anonymous):

I don't understand this statement on this website . Please help, this is about half-angle identities and +/-.

OpenStudy (anonymous):

The statement in question is this: http://bit.ly/AuIXO0

OpenStudy (nottim):

Try Khan's Academy?

OpenStudy (anonymous):

I tried that, but there was nothing

OpenStudy (nottim):

no txtbook?

OpenStudy (anonymous):

No Txtbook doesn't explain. Im confused about the +/- sign for the final answer. I do everything right except that.

OpenStudy (phi):

They ask about Find the possible values of tan(1/2)theta when tantheta = 7/24 This means \[ tan(\theta) = \frac{7}{24} \] or \[ tan(\theta) = \frac{-7}{-24} \] The first angle is in the 1st quadrant and the half angle will be in the first quadrant the second angle is in the third quadrant, and half of it is in the 2nd quadrant. (So its tan will be negative)